Moonscape has just completed an initial public offering. The firm sold 4 million shares at an offer price of $12 per share. The underwriting spread was $0.40 a share. The price of the stock closed at $18.00 per share at the end of the first day of trading. The firm incurred $400,000 in legal, administrative, and other costs. What were flotation costs as a fraction of funds raised? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
Answer #1

Fund raised = shares issued*offer price

Fund raised = 4*12 = 48

flotation cost as percentage of fund raised = total flotation cost / fund raised = 34/48 = .7083 = 70.83%

total flotation cost = Underwriting cost + Under pricing + other, legal, administrative and other cost

total flotation cost =(4*.4)+ 4*(18-10)+.4 = 1.6+32+.4 = 34

fund raised = 48